# Enabling Ensemble Learning for Heterogeneous Large Language Models with Deep Parallel Collaboration

### v1

@@ Abstract
Large language models (LLMs) have shown complementary strengths in various tasks and instances, motivating the research of ensembling LLMs to push the frontier leveraging the wisdom of the crowd. Existing work achieves this objective via training the extra reward model or fusion model to select or fuse all candidate answers. However, these methods pose a great challenge to the generalizability of the trained models. Besides, existing methods use the textual responses as communication media, ignoring the rich information in the inner representations of neural networks. Therefore, we propose a training-free ensemble framework DeePEn, averaging the probability distributions outputted by different LLMs. A key challenge in this paradigm is the vocabulary discrepancy between heterogeneous LLMs, which hinders the operation of probability distribution averaging. To address this challenge, DeePEn maps the probability distribution of each model from the probability space to a universe relative space based on the relative representation theory, and performs aggregation. Then, the result of aggregation is mapped back to the probability space of one LLM via a search-based inverse transformation to determine the generated token. We conduct experiments on the ensemble of various LLMs of 6B to 70B. Experimental results show that DeePEn achieves consistent improvements across six popular benchmarks involving subject examination, reasoning and knowledge-QA, proving the effectiveness of our approach [^1] Our code will be released soon.. 🖂 means corresponding author.
{CJK}
UTF8gbsn
Enabling Ensemble Learning for Heterogeneous Large Language Models with Deep Parallel Collaboration
Yichong Huang, Xiaocheng ${}^{{\dagger}{\ddagger}\textrm{\Letter}}$, Baohang Li, Yang Xiang, Hui Wang Ting Liu, Bing Qin†‡ Harbin Institute of Technology    Peng Cheng Laboratory {ychuang,xcfeng,baohangli,tliu,qinb}@ir.hit.edu.cn {xiangy,wangh06}@ir.hit.edu.cn
@@ 1 Introduction
With the scaling of model capacities and data volumes, generative large language models (LLMs) have shown impressive language understanding and generation abilities, shedding light for artificial general intelligence Zhao et al. (2023); OpenAI (2023); Jiang et al. (2023a); Touvron et al. (2023). Due to diversities of data sources, model architectures, and training recipes, LLMs have different strengths and weaknesses in various tasks and cases. Therefore, recent research has explored the ensemble of LLMs to exploit the complementary potential (Jiang et al., 2023b; Lu et al., 2023).
Existing work employs a two-stage process for LLM ensemble, where each LLM first generates candidate answers independently, and then the final answer is derived by either selecting or fusing these candidates, corresponding to selection-based and fusion-based ensemble approaches. Selection-based ensemble leverages an extra trained reward model to score all candidates and select the best one (Jiang et al., 2023b; Wang et al., 2024; Shnitzer et al., 2024; Lu et al., 2023), which fails to achieve collaboration between LLMs. Fusion-based ensemble trains a fusion model to generate a better answer according to all candidates (Jiang et al., 2023b) or adopt the LLM to perform fusion (Du et al., 2023). However, these methods either pose a great challenge to the generalizability of the reward model and the fusion model or incur a non-trivial computation cost and inference latency for communication. Besides, existing research enables collaboration via conveying the textual responses between LLMs while ignoring the rich information (e.g., confidence) in the inner representations.
@@
One ideal solution to this dilemma is prediction fusion (Bisk et al., 2020; Garmash and Monz, 2016). For LLM ensemble, prediction fusion works at each decoding step, averaging the probability distributions from different LLMs to determine the generated token. This paradigm is not only training-free, making it more general, but also leverages the inner representations (i.e., probability distributions) as communication media between LLMs. An essential assumption for prediction fusion is that the ensemble models share a common vocabulary. However, this assumption often does not hold due to the heterogeneity of LLMs caused by different vocabulary sizes and tokenizer models.
@@
In this work, we tackle this key challenge via drawing upon the cross-model invariance of relative representation, which represents each token using the embedding similarities of this token to a set of anchor tokens (Moschella et al., 2023). Specifically, we propose an ensemble framework DeePEn (Deep Parallel Ensemble), which enables prediction fusion for heterogeneous LLMs. DeePEn transforms the probability distribution from the heterogeneous probability space to a homogeneous relative space, using a matrix formed by the relative representation of all tokens. We refer to this matrix as a relative representation matrix. Next, DeePEn aggregates the relative representations of all probability distributions in the relative space, to obtain the representation containing the comprehensive decision on the tokens to generate. Finally, the result of aggregation are transformed back to the probability space of one of the ensemble LLMs using a search-based inverse transformation to determine the generated token.
Although our approach can apply on the ensemble of any number of LLMs, we mainly prove its effectiveness on the 2-model and 3-model ensemble. Our experiments involve various LLMs of 6B to 70B, and consider dense models and the MoE model Mixtral. Experimental results demonstrate that DeePEn achieves consistent improvement on six widely-used benchmarks.
@@ 2 Theoretical Analysis
In this section, we first introduce relative representation and then illustrate the theoretical support for our methodology.
Figure 1: Visualizations for relative representations between homogeneous models and between and heterogeneous models. PCA is applied only for visualization. Red block indicate the representation of tokens that only appear in the vocabulary of Mistral. Relative representation consistency is obtained by calculating the cosine similarity between the relative representations of the same token in different models.
Figure 2: Overview of DeePEn.
@@ 2.1 Relative Representation
Previous study discovers that despite the misalignment between latent spaces of different neural networks, the embedding similarity between samples do not change across models (Moschella et al., 2023; He and Ozay, 2022; Park et al., 2019). Specifically, Moschella et al. (2023) propose relative representation, which represents each sample $x^{(i)}$ by the embedding similarities to a set of anchor samples $\mathbb{A}$. Formally:
@@
where $e_{(*)}$ denotes the embedding of samples, also is absolute representation. Note that $x^{(i)}$ and anchors $\mathbb{A}$ are identically distributed.
It is empirically discovered that relative representation possesses cross-model invariance, i.e., the relative representation of the same sample keeps invariant across different models, which lays the theoretical foundation for our work to ensemble heterogeneous LLMs.
@@ 2.2 Theoretical Support for DeePEn
Averaging probability distribution (i.e., prediction fusion) is an effective method to achieve the coordination between various neural networks. The underlying mechanism is to interpolate the output semantics (i.e., the semantic that model intends to express) in the probability space. However, vocabulary discrepancy isolates these target semantics in semantic spaces with different basis vectors, hindering the interpolation. To tackle this challenge, we aim to enable the cross-model alignment for target semantics, i.e., find a transformation to transform the target semantic into a universal space. To this effect, we propose to represent the target semantic with the convex combination of relative representations of all tokens where the weight is the probability assigned to the token.
Definition of target semantics in relative space.
Formally, given the absolute representation of the output semantic $\mathbf{p}^{(t)}$ and the relative representation matrix $R\in\mathbb{R}^{|V|\times|A|}$ where $V$ is the vocabulary and $A\subseteq V$ is the anchor token set. The $i$-th row of $R$ is the relative representation of word $w^{(i)}$:
@@
and the relative representation of the target semantic $\mathbf{p}^{(t)}$ is defined as:
$\mathbf{r}^{(t)}=\mathbf{p}^{(t)}\cdot R$ (3)
Model-invariance of relative representation of target semantic.
Next, we illustrate why this representation scheme could align the output semantics isolated in heterogeneous absolute spaces. First, considering two LLMs $\theta_{A}$ and $\theta_{B}$ with the same vocabulary (e.g., LLaMA2-7B and LLaMA2-13B). When expressing the same target semantic, these models output the same probability distribution (i.e., absolute representation) $\mathbf{p}^{(t)}_{A}$ and $\mathbf{p}^{(t)}_{B}$. Besides, they have the same (highly similar in practice) relative representation matrix due the vocabulary consistency and cross-model invariance of relative representation. Therefore, the induced relative representations of output semantics are also the same:
$\mathbf{r}^{(t)}_{A}$ $=\mathbf{p}^{(t)}_{A}\cdot R_{A}=\mathbf{p}^{(t)}_{B}\cdot R_{B}=\mathbf{r}^{(t)}_{B}.$ (4)
Then, let’s consider a language model $\theta_{C}$ with a different vocabulary (e.g., Mistral). Based on the fact that different LLMs share mass tokens in their vocabularies, the vocabulary of model $\theta_{C}$ is identical to adding and removing partial tokens to the vocabulary of $\theta_{B}$, which leads to $\mathbf{p}^{(t)}_{B}\neq\mathbf{p}^{(t)}_{C}$ and $R_{B}\neq R_{C}$. However, in our study, we discover that this change to vocabularies has not incurred significant influence on the relative representation of the unchanged tokens (i.e., the common tokens between $\theta_{B}$ and $\theta_{C}$), as shown in Fig. 1. Therefore, we assume that the local change in the vocabulary could hardly influence the relative space. This hypothesis is strict but reasonable as the semantic does not change with the tokenizer intuitively.
@@ 3.1 Overview
Given $N$ models to ensemble, DeePEn first constructs their transformation matrices (i.e., relative transformations) mapping the probability distributions from the heterogeneous absolute spaces into the relative space (§3.2). At each decoding step, all models perform the prediction and output $N$ probability distributions. These distributions are transformed into the relative space and aggregated (§3.3). Finally, the result of aggregation is transformed back into the absolute space of one of the ensembling models, and is used to determine the generated token (§3.4).
@@ 3.2 Construction of Relative Transformation
The choice of anchor tokens is crucial for the relative representation capability. Previous research discovers that the capability improves as the number of anchor words increases Moschella et al. (2023). Therefore, we employ the full set of common words between LLMs as the anchor words. It is also empirically proved that this method achieves more stable performance on downstream tasks.
@@ 3.3 Aggregation in Relative Space
At each decoding step, once each model $\theta_{i}$ outputs the probability distribution $\textbf{p}_{i}$, DeePEn transforms $\textbf{p}_{i}$ into the relative representation $\textbf{r}_{i}$ using the normalized relative representation matrix:
$\mathbf{r}_{i}=\textbf{p}_{i}\cdot\hat{R}_{i},$ (6)
@@
$\overline{\mathbf{r}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{r}_{i}.$ (7)
Note that we adopt the most simple method of aggregation (i.e., averaging) to make our framework more general in this work. But it could be integrated with other sophistic technologies of output fusion (e.g., stacking (Wolpert, 1992))
@@ 3.4 Inverse Transformation of Relative Representations
To decide the emitted token according to the result of aggregation, DeePEn aims to transform it from the relative space back to the absolute space of one of ensembling models (main model). In practice, we select the model with the best performance on the validation set as the main model, which often achieves better ensemble performance. To enable this inverse transformation, we adopt a research strategy, which finds out the absolute representation that the relative representation of which is identical to the aggregated relative representation. This search problem is formulated as:
$\overline{\textbf{p}}_{(i)}=\mathop{\arg\min}\limits_{\textbf{p}_{(i)}^{\prime}\in\ \mathbb{P}_{(i)}}\ell(\textbf{p}_{(i)}^{\prime}\times\hat{R},\ \overline{\mathbf{r}}),$ (8)
where $\mathbb{P}_{(i)}$ denotes the absolute space of model $\theta_{i}$, and $\ell(\cdot)$ is the loss function to measure the distance between two relative representations. In this work, we adopt the KL-divergence since for its stability in our experiments.
This search is conducted under the guidance of the gradient of the divergence between relative representations with respect to the absolute representation. Specifically, the start point is initialized with the main model’s original absolute representation $\textbf{p}_{(i)}$, and use the gradient of the loss in Eq.8 with respect to $\textbf{P}_{(i)}^{\prime}$:
$\textbf{p}_{(i)}^{\prime}=\textbf{p}_{(i)}^{\prime}-\eta\times\frac{\partial\ell}{\partial\textsc{p}_{(i)}^{\prime}},$ (9)
where $\eta$ is an important hyperparameter named as the relative ensemble learning rate. This search process is iterated for $T=5$ steps, where $T$ is named as the number of relative ensemble learning steps.
Finally, we use the updated absolute representation $\textbf{p}_{(i)}^{\prime}$ to determine the emitted token.
@@ 4.1 Experimental Setup
Although DeePEn is a task-agnostic approach, which could apply on any NLP task or even beyond NLP, we mainly conduct experiments on six benchmarks, which could be categorized as three kinds of tasks:
  • Comprehensive Examination: (1) MMLU (5-shot) Hendrycks et al. (2021), which covers 57 subjects that humans learn across STEM, the humanities, and the social sciences and ranges in difficulty from an elementary level to an advanced professional level, and (2) ARC-C (0-shot) Clark et al. (2018), which consists of a collection of natural science questions authored for standardized tests.
  • Reasoning Capabilities: (1) GSM8K Cobbe et al. (2021) (4-shot), which is a dataset of high quality problems at the grade school math level, and (2) PIQA (Bisk et al., 2020) (0-shot), which is a commonsense reasoning dataset.
  • Knowledge Capacities: (1) TriviaQA (5 shot) Joshi et al. (2017), which consists of Trivia enthusiast authored question-answer pairs, and (2) NaturalQuestions (5-shot) Kwiatkowski et al. (2019), which is a QA corpus consists of queries issued to the Google search engine.
Models MMLU ARC-C
dev test dev test
@@
Yi-6B 61.19 63.24 72.72 73.33
Mistral-7B 60.80 62.07 73.88 74.10
SkyWork-13B 58.65 61.21 67.08 66.50
@@
DeePEn 63.61 65.69 77.73 75.89
$\Delta$ +2.42 +2.45 +3.85 +1.79
Top-3 Ensemble
PairRanker 63.77 70.85
GenFuser 37.59 58.21
@@
$\Delta$ +2.06 +2.01 +4.21 +2.99
Table 1: Results on comprehensive examination.
Models GSM8K PIQA
dev test dev test
Individual Models
LLaMA2-70B 68.67 65.73 70.54 71.27
Mixtral-8$\times$7B 66.67 63.84 70.90 71.88
@@
$\Delta$ +1.00 +1.60 +2.64 +3.22
Table 2: Results on reasoning tasks.
Evaluation.
For all benchmarks, we follow the test scripts of OpenCompass [^2] https://opencompass.org.cn/. Specifically, on the multiple-choice tasks (MMLU, ARC-C, and PIQA), the option with the highest likelihood is selected to calculate the accuracy. On the free-form generation tasks (GSM8K, TrivialQA and NaturalQuestions), we calculate the exact match (EM) accuracy.
Individual models.
Based on the key principles of ensemble learning Sagi and Rokach (2018), the synergy is created only in cases where individual models possess comparable performance to each other. Therefore, on each task type, we select three LLMs that have achieved promising results on the OpenCompass leaderboard and whose performance is as close to each other as possible. On the challenge reasoning task, we conduct experiments for the ensemble of two large-scale LLMs: LLaMA2-70B and Mixtral-8$\times$7B.
Hyperparameters.
In this work, we select all of the common tokens between LLMs as the anchor tokens to build the relative spaces, i.e., $A=C$. In the decoding of relative representation, we search the optimal relative learning rate (Eq. 9) from 0.0 to 0.3 with an interval of 0.05. We set the number of relative ensemble learning steps $T=5$.
Comparative methods.
We also compare our ensemble framework DeePEn with the selection-based ensemble method PairRanker, which is a reward model to score each response of LLMs and the fusion-based ensemble method GenFuser, which is a generative model to fuse multiple candidate responses. Both models are from LLM-Blender framework Jiang et al. (2023b) and are trained on the constructed instruction tuning dataset MixInstruct.
Models TriviaQA NQ
dev test dev test
Individual Models
LLaMA2-13B 72.74 73.57 21.56 28.67
Mistral-7B 70.47 72.58 19.94 27.67
InternLM-20B 64.11 65.90 19.50 26.09
Top-2 Ensemble
DeePEn 74.79 75.13 21.81 30.17
$\Delta$ +2.05 +1.56 +0.25 +1.50
Top-3 Ensemble
PairRanker 70.27 26.9
GenFuser 1.87 0.17

### v2

# Ensemble Learning for Heterogeneous Large Language Models with Deep Parallel Collaboration

@@ Abstract
Large language models (LLMs) exhibit complementary strengths in various tasks, motivating the research of LLM ensembling. However, existing work focuses on training an extra reward model or fusion model to select or combine all candidate answers, posing a great challenge to the generalization on unseen data distributions. Besides, prior methods use textual responses as communication media, ignoring the valuable information in the internal representations. In this work, we propose a training-free ensemble framework DeePEn, fusing the informative probability distributions yielded by different LLMs at each decoding step. Unfortunately, the vocabulary discrepancy between heterogeneous LLMs directly makes averaging the distributions unfeasible due to the token misalignment. To address this challenge, DeePEn maps the probability distribution of each model from its own probability space to a universal relative space based on the relative representation theory, and performs aggregation. Next, we devise a search-based inverse transformation to transform the aggregated result back to the probability space of one of the ensembling LLMs (main model), in order to determine the next token. We conduct extensive experiments on ensembles of different number of LLMs, ensembles of LLMs with different architectures, and ensembles between the LLM and the specialist model. Experimental results show that (i) DeePEn achieves consistent improvements across six benchmarks covering subject examination, reasoning, and knowledge, (ii) a well-performing specialist model can benefit from a less effective LLM through distribution fusion, and (iii) DeePEn has complementary strengths with other ensemble methods such as voting [^1] Our code is available at: https://github.com/OrangeInSouth/DeePEn.
@@ 1 Introduction
With the scaling of model capacities and data volumes, generative large language models (LLMs) have shown impressive language understanding and generation abilities, shedding light for artificial general intelligence [35, 22, 13, 28]. Due to diversities of data sources, model architectures, and training recipes, LLMs have different strengths and weaknesses in various tasks and cases. Therefore, recent research has explored the ensemble of LLMs to exploit the complementary potential [15, 19].
Existing methods can be categorized into selection-based and fusion-based ensembling. Selection-based ensembling selects the best candidate answer from all individual LLMs’ answers using an additionally trained reward model [15, 31, 25, 19]. Fusion-based ensembling combines all candidate answers using a trained fusion model [15]. However, these approaches inevitably face significant challenges in generalizing to unseen data distributions and base models. Besides, prior methods enable collaboration via conveying the textual responses between LLMs while ignoring the rich information (e.g., confidence and alternative answers) in the internal representations.
@@
An ideal solution to this issue is to apply the well-established technology of prediction fusion. [36, 24, 7, 10]. For LLM ensemble, prediction fusion works at each decoding step, averaging the probability distributions from different LLMs to determine the next token. It could not only directly apply to the ensemble of any LLMs without extra parameter training, making it more general, but leverages the informative internal representations (i.e., probability distributions) as communication media. Unfortunately, the vocabulary discrepancy between different LLMs makes it unfeasible to average the distributions due to token misalignment.
@@
In this work, we tackle this key challenge by drawing upon the cross-model invariance of relative representation, which represents each token using the embedding similarities of this token to a set of anchor tokens [21]. Specifically, we propose an ensemble framework DeePEn (Deep Parallel Ensemble), enabling distribution fusion for heterogeneous LLMs. DeePEn transforms the probability distribution from the heterogeneous probability space to a homogeneous relative space, using a matrix formed by the relative representation of all tokens. Next, DeePEn aggregates the relative representations of all probability distributions in the relative space, coordinating the decision on the next token. Finally, the result of aggregation is transformed back to the probability space of the main model using a search-based inverse transformation to determine the next token.
We conduct extensive experiments ranging from 2-model to 9-model ensembles, covering ensembles of models with parameters ranging from 6B to 70B, ensembles of dense and sparse models, and the ensemble of LLMs with specialist models. Experimental results on six widely-used benchmarks demonstrate that compared to baselines, DeePEn achieves consistent improvements across all benchmarks. It is also discovered that DeePEn has complementary strengths when combined with other ensemble methods.
@@ 2 Theoretical Analysis
We first introduce relative representation and then illustrate the theoretical support for our method.
@@ 2.1 Relative Representation
Previous study discovers that despite the misalignment between latent spaces of different neural networks, the embedding similarity between samples do not change across models [21, 11, 23]. Specifically, Moschella et al. [21] propose relative representation, which represents each sample $x^{(i)}$ by the embedding similarities to a set of anchor samples $\mathbb{A}$ ($x^{(i)}$ and $\mathbb{A}$ are identically distributed):
@@
where $e_{(*)}$ denotes the embedding of samples, also is absolute representation.
It is empirically evidenced that relative representations possess cross-model invariance, i.e., the relative representation of the same sample keeps invariant across different models, which lays the theoretical foundation for our work to fuse heterogeneous probability distributions.
@@ 2.2 Theoretical Support for DeePEn
Figure 1: Visualizations for relative representations between models with the same vocabulary and between models with different vocabularies. PCA and K-means clustering are applied only for visualization. The red block indicates the representation of tokens that only appear in Mistral’s vocabulary. Relative representation consistency is obtained by calculating the cosine similarity between the relative representations of the same token in different models.
Average probability distribution has been widely evidenced to effectively improve the predictive performance in the filed of image and text [2, 10]. For generative language models, as we understand, the underlying mechanism is to interpolate different output semantics represented by the probability distributions. However, for LLM ensemble, vocabulary discrepancy isolates these output semantics in semantic spaces with different basis vectors, making the interpolation infeasible. To tackle this challenge, we aim to enable the cross-model alignment for output semantics, i.e., find a transformation to map the output semantics into a universal space. To this effect, we propose to represent the output semantics with the convex combination of relative representations of all tokens where the weight is the probability assigned to the token.
Definition of output semantics in relative space.
Formally, given the absolute representation of the output semantics $\mathbf{p}$ and the relative representation matrix $R\in\mathbb{R}^{|V|\times|A|}$ where $V$ is the vocabulary and $A\subseteq V$ is the anchor token set. The $i$-th row of $R$ is the relative representation of word $w^{(i)}$:
@@
and the relative representation of the output semantics $\mathbf{p}$ is defined as: $\mathbf{r}=\mathbf{p}\cdot R$.
Model-invariance of relative representation of output semantic.
Next, we illustrate why this representation scheme could align the output semantics isolated in heterogeneous absolute spaces. First, considering two LLMs $\theta_{A}$ and $\theta_{B}$ with the same vocabulary (e.g., LLaMA2-7B and LLaMA2-13B). When expressing the same output semantic, these models output the same probability distribution (i.e., absolute representation) $\mathbf{p}_{A}$ and $\mathbf{p}_{B}$. Besides, they have the same (highly similar in practice) relative representation matrix due the vocabulary consistency and cross-model invariance of relative representation. Therefore, the relative representations of output semantics are also identical:
$\mathbf{r}_{A}$ $=\mathbf{p}_{A}\cdot R_{A}=\mathbf{p}_{B}\cdot R_{B}=\mathbf{r}_{B}.$ (3)
Then, let’s consider a language model $\theta_{C}$ with a different vocabulary (e.g., Mistral). Based on the fact that different LLMs typically share mass tokens in their vocabularies (§A), the vocabulary of model $\theta_{C}$ is identical to adding and removing partial tokens to the vocabulary of $\theta_{B}$, which leads to $\mathbf{p}_{B}\ncong\mathbf{p}_{C}$ and $R_{B}\ncong R_{C}$. However, in our study, we discover that this change to the vocabulary has not incurred significant influence on the relative representation of the unchanged tokens (i.e., the common tokens between $\theta_{B}$ and $\theta_{C}$), as shown in Fig. 1. Therefore, we make the reasonable assumption that the local change in the vocabulary could hardly influence the relative space.
@@ 3.1 Overview
We illustrate the process of DeePEn in Fig. 2. Given $N$ models to ensemble, DeePEn first constructs their transformation matrices (i.e., relative representation matrices) mapping the probability distributions from the heterogeneous absolute spaces into the relative space (§3.2). At each decoding step, all models perform prediction and output $N$ probability distributions. These distributions are mapped into the relative space and aggregated (§3.3). Finally, the aggregation result is transformed back into the absolute space of the main model, in order to determine the next token (§3.4).
Figure 2: Overview of DeePEn. The relative representation matrix of each LLM is directly derived by calculating the embedding similarities between each token with the anchor tokens.
@@ 3.2 Construction of Relative Transformation
The choice of anchor tokens is crucial for the relative representation capability. Previous research discovers that the capability improves as the number of anchor words increases [21]. Therefore, we employ the full set of common words between LLMs as the anchor words. It is also empirically proved that this method performs more stablely on downstream tasks (§5.2).
@@ 3.3 Aggregation in Relative Space
At each decoding step, once each model $\theta_{i}$ outputs the probability distribution $\textbf{p}_{i}$, DeePEn transforms $\textbf{p}_{i}$ into the relative representation $\textbf{r}_{i}$ using the normalized relative representation matrix: $\mathbf{r}_{i}=\textbf{p}_{i}\cdot\hat{R}_{i}$,
@@
$\overline{\mathbf{r}}=\sum_{i=1}^{N}\alpha_{i}\times\mathbf{r}_{i},$ (5)
where $\alpha_{i}$ is the collaboration weight of model $\theta_{i}$3.5).
@@ 3.4 Inverse Transformation of Relative Representations
To decide the next token according to the aggregated relative representation, DeePEn aims to transform it from the relative space back to the absolute space of the main model, which is empirically selected with the best-performing model on the development set. To enable this inverse transformation, we adopt a search-based strategy, finding out the absolute representation whose relative representation is identical to the aggregated relative representation. This search problem is formulated as:
$\overline{\textbf{p}}_{i}=\mathop{\arg\min}\limits_{\textbf{p}_{i}\in\ \mathbb{P}_{i}}\ell(\textbf{p}_{i}\times\hat{R},\ \overline{\mathbf{r}}),$ (6)
where $\mathbb{P}_{i}$ denotes the absolute space of model $\theta_{i}$, and $\ell(\cdot)$ is the loss function to measure the distance between relative representations. In this work, we adopt the KL-divergence due to its convergence.
This search is iteratively conducted under the guidance of the gradient of the loss in Eq.6 with respect to the absolute representation $\textbf{p}_{i}$. Specifically, we initialize the start point of searching $\textbf{p}^{(0)}_{i}$ with the main model’s original absolute representation, and update it as:
$\textbf{p}^{(t+1)}_{i}=\textbf{p}^{(t)}_{i}-\eta\times\frac{\partial\ell}{\partial\textbf{p}^{(t)}_{i}},t\in[0,T]$ (7)
where $\eta$ is an important hyperparameter named the relative ensemble learning rate, and $T$ is the iterations number named relative ensemble learning steps. Finally, we use the updated absolute representation $\textbf{p}^{(T)}_{i}$ to determine the emitted token.
3.5 Collaboration Schemes
DeePEn aggregates the output distributions of individual models via performing weighted averaging on their relative representations (Eq. 5). As our work focus on enabling the distribution fusion of heterogeneous LLMs instead of finding the optimal collaboration weights, we follow the most common practice to uniformly aggregate the distributions ($\alpha=1/N$, $N$ is the number of models), which is named DeePEn-Avg. Besides, we also adopt a simple and effective method of deducing weights, DeePEn-Adapt, which heuristically sets a larger value to the model with a better performance on the development set: $\alpha_{i}=s_{i}/\sum_{j}s_{j}$, where $s_{i}=Acc(\theta_{i},\mathcal{D}^{dev})-\epsilon$, $Acc(\cdot,\cdot)$ indicates the average accuracy of model $\theta_{i}$ on the development set, and $\epsilon$ indicates the chance level on the evaluation task. Specifically, $\epsilon=0$ on the free-form generation tasks and $\epsilon=1/K$ on the $K$-choice tasks.
@@ 4.1 Experimental Setup
We mainly conduct experiments on six benchmarks, which can be categorized into:
  • Comprehensive Examination: (1) MMLU (5-shot) [12], which covers 57 subjects that humans learn, and (2) ARC-C (0-shot) [5], collected from standardized natural science tests.
  • Reasoning Capabilities: (1) GSM8K [6] (4-shot), which is a dataset of high quality problems at the grade school math level, and (2) PIQA [3] (0-shot), which is a commonsense reasoning dataset.
  • Knowledge Capacities: (1) TriviaQA (5-shot) [16], collected by Trivia enthusiast authored, and (2) NQ (5-shot) [18], which is a QA corpus consists of queries issued to the Google search engine.
Evaluation.
For all benchmarks, we follow the test scripts of OpenCompass leaderboard. Specifically, on the multiple-choice tasks (MMLU, ARC-C, and PIQA), the option with the highest likelihood is selected to calculate the accuracy. On the free-form generation tasks (GSM8K, TriviaQA and NQ), we calculate the exact match (EM) accuracy.
Individual models.
As ensemble learning typically works on models with comparable performance [24, 34], we select six well-performing LLMs whose performance are closely matched: LLaMA-2-13B [29], Mistral-7B-v0.1 [13], InternLM-20B [26], Yi-6B [1], Skywork-13B-base [32], and Tigerbot-13b-base-v2 [4]. To achieve better ensemble performance, we conduct experiments on the ensemble of the top-2 models and the top-4 models for each benchmark. Besides, we also consider ensembling various number of models (§4.3) and ensembling more diverse models (§5.1).
Hyperparameters.
[…]
Comparative methods.
We compare DeePEn with (1) MinED [30, 9], which maps the probability distributions of heterogeneous LLMs to the distribution of the main model via aligning tokens in different vocabularies with edit distance, and (2) LLM-Blender [15], which comprises a reward model PairRanker to score each response of LLMs and a fusion model GenFuser to fuse candidate responses. In this work, we we only adopt the PairRanker since GenFuser suffers from serious over-generation under our training-free setting. In the ensemble of more than two models, we introduce two additional ensemble methods: (3) Voting, which selects the choice favored by most models on the tasks with outputs limited to a fixed set, and (4) MBR [8, 17], which selects the answer with the highest textual similarity to other candidate answers. The implementation details of baselines are illustrated in §B.
Models Examination Reasoning Knowledge
MMLU ARC-C GSM8K PIQA TriviaQA NQ
@@
LLaMA2-13B 55.07 59.32 29.80 59.68 74.32 28.67
InternLM-20B 59.94 75.81 53.83 64.78 66.88 26.09
Skywork-13B 61.16 66.50 53.90 74.04 58.65 19.75
Tigerbot-13B 51.95 57.44 48.82 68.28 66.22 22.71
Mistral-7B 62.13 73.33 47.50 65.61 73.18 27.62
Yi-6B 63.25 73.33 37.91 76.15 59.02 18.98
@@
LLM-Blender 63.85 (+0.60) 75.73 (- 0.08) 54.89 (+0.99) 78.31 (+2.16) 74.10 (- 0.22) 28.61 (- 0.06)
MinED 65.04 (+1.79) 77.35 (+1.54) 18.50 (-35.40) 78.98 (+2.83) 72.30 (- 2.02) 28.45 (- 0.22)
@@
Top-4 Ensemble
LLM-Blender 61.44 (- 1.81) 71.03 (- 4.78) 43.37(-10.53) 71.16 (- 4.99) 67.87 (- 6.45) 24.18 (- 4.49)
Voting 64.88 (+1.63) 78.41 (+2.60) 63.15 (+9.25) 76.82 (+0.67)
MBR 62.09 (+8.26) 74.32 (+0.00) 30.28 (+1.61)
MinED 65.61 (+2.36) 78.68 (+2.87) 56.56 (+2.66) 77.87 (+1.72) 71.62 (- 2.70) 29.50 (+0.83)
@@
DeePEn-Adapt 65.25 (+2.00) 79.15 (+3.34) 56.25 (+2.35) 78.59 (+2.44) 75.76 (+1.44) 31.77 (+3.10)
+Voting/MBR 65.40 (+2.15) 79.44 (+3.63) 65.25 (+11.35) 77.37 (+1.22) 75.65 (+1.33) 32.11 (+3.44)
Table 1: Main results. The best individual model is highlighted in red, and the best ensemble method is highlighted in green , except for the results of the combined method (i.e., the last row). The top-4 models on each benchmark are underlined. ‘—’ indicates that the method does not apply to the task.
4.2 Main Results
The main results are shown in Tab. 4.1, from which we have drawn the following observations:
(1) DeePEn achieves consistent improvements over the individual models.
These results prove that our DeePEn successfully enables collaboration between heterogeneous LLMs via aggregating their probability distributions in the relative space. Specifically, DeePEn-Avg achieves improvements of +1.43(MMLU)$\sim$+2.72(PIQA) on the ensemble of top-2 models, and +1.00(PIQA)$\sim$+2.89(ARC-C) on the ensemble of top-4 model. DeePEn-Adapt gains improvements of +1.44(TriviaQA)$\sim$+3.34(ARC-C) on the ensemble of top-4 models.
(2) DeePEn shows better stability than baselines.
As shown, LLM-Blender struggles to achieve improvements under the training-free setting. MinED shows unstable performance across different benchmarks. For example, MinED leads to performance drops of -35.40 on the GSM8K benchmark under the top-2 models ensemble setting and -2.70 on the TriviaQA, indicating the limitation of using textual similarity to align tokens in heterogeneous vocabularies. Through case studies, it is revealed that this method of aligning tokens with edit distance disturbs the decoding and produces incomplete words (demonstrated in §10). Instead, DeePEn-Avg achieves consistent improvements and surpasses all baselines in 7/12 settings.
(3) DeePEn has complementary strengths with other ensemble methods.
Voting achieves a significant improvement on the mathematical reasoning GSM8K, showing the effectiveness of reasoning with multiple paths. To evidence the complementary strength of DeePEn with Voting, we combine both methods. On the TriviaQA and NQ, Voting is replaced with MBR. As shown that the combination of both methods gains a further improvement over Voting (63.15$\rightarrow$65.25).
(4) Collaboration with more worse-performing LLMs is a double-edged sword.
The ensemble performance of DeePEn-Avg with top-4 models surpasses that with top-2 models on 4 benchmarks, but falls short on 2 benchmarks. This is reasonable because incorporating the 3rd and 4th ranked LLMs enhances complementary strengths but also causes the interference with the top-2 models.
4.3 Results on Different Numbers of Models
Figure 3: Test set results of ensemble learning on various number of models. Individual models are arranged in descending order of their performance on the development set, and sequentially incorporated into the ensemble. $\hat{\Delta}$ indicates the largest improvement achieved by DeePEn.
Next, we illustrate the effectiveness of DeePEn on the ensemble of more models on the MMLU, PIQA, and NQ. We add Nanbeige-13B into the ensemble on all three benchmarks, and add LLaMA2-70B and Mixtral-8$\times$7B on the PIQA due to their comparable performance. As illustrated in Fig. 3, the ensemble performance increases first and then decreases with the joining of more models in descending order of performance. And the ensemble performance peaks in the top-4 or top-5 models across three benchmarks.
5 Analysis
To deeply understand DeePEn, we first evaluate its performance on the ensemble learning of model sets with diverse architectures, abilities, and performance gaps. Next, we conduct a series of analyses on the reverse transformation process of relative representations.
5.1 Results of Ensembling Diverse Models
Model GSM8K PIQA
LLaMA2-70B (Dense) 63.84 71.27
Mixtral-8$\times$7B (Sparse) 65.73 71.88